Calculate the pH of a 1.0 M solution of methylamine (Kb = 4.38 x 104
).
1
Expert's answer
2020-05-25T14:13:26-0400

Solution:

Since methylamine (CH3NH2) is a weak base, the major species in solution are CH3NH2 and H2O.

Both are bases; however, since water can be neglected as a source of OH-, the dominant equilibrium is:

CH3NH2(aq) + H2O(l) = CH3NH3+(aq) + OH-(aq)

and



The concentrations are as follows:



[OH-] = [CH3NH3+] = x

[CH3NH2] = 1.0 - x

Then,




Assume that 1 >> x, then:

(1.0 - x) = 1.0

Kb = x2 / 1.0

x2 = Kb

x = (Kb)0.5 = (4.38×10-4)0.5 = 0.0209

x= [OH-] = 0.0209 M


We can convert between [OH-] and pOH using the following equations:

pOH = - log[OH-]

pOH = - log(0.0209) = 1.68

pOH = 1.68


For any aqueous solution at 25C:

pH + pOH = 14

pH = 14 - pOH = 14 - 1.68 = 12.32

pH = 12.32


Answer: pH = 12.32